what is that answer to 2x+3y=12 4x=10+y when solving by elimination

Answer by TimothyLamb(4379)

---
2x + 3y = 12
4x = 10 + y
---
put the system of linear equations into standard form
---
2x + 3y = 12
4x - y = 10
---
copy and paste the above standard form linear equations in to this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer by matrix method:
x= 3
y= 2
---
---
now by elimination:
---
2x + 3y = 12
12x - 3y = 30
---
add:
14x = 42
x = 3
2*3 + 3y = 12
6 + 3y = 12
3y = 6
y = 2
---
answer by elimination:
x= 3
y= 2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php